OpenStudy (anonymous):

Solve each system algebracally 10x+3y=19 y=2x+5

OpenStudy (anonymous):

you know what y is from equation 2 so put the value of y, (2x + 5), into the y for the top equation and solve for x. Tell me where you get stuck (if anywhere)

OpenStudy (anonymous):

so 10x + 3(2x + 5) = 19 solve for x

OpenStudy (anonymous):

x=1? @JamesWolf

OpenStudy (anonymous):

erm lemmi see 10x + 6x + 15 = 19 10x + 6x = 4 x(10+6) = 4 x = 4/16 x = 1/4

OpenStudy (anonymous):

you agree?

OpenStudy (anonymous):

yes i do i didn't add like terms

OpenStudy (anonymous):

where does 15 go?

OpenStudy (anonymous):

so 10x + 6x + 15 = 19 take 15 from both sides 10x + 6x + 15 - 15 = 19 - 15 10x + 6x = 4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@JamesWolf